OpenStudy (anonymous):

Help! Solve for x. 2^x-3=128

OpenStudy (vera_ewing):

Do you know what the first step is?

OpenStudy (anonymous):

I don't

OpenStudy (campbell_st):

well you need to rewrite 128 as a power of 2 \[128 = 2^7\] so the problem becomes \[2^{x -3} = 2^7\] so equate the powers and solve for x x - 3 = 7

OpenStudy (anonymous):

10?

OpenStudy (campbell_st):

questions of this type, you should re-write them with the same base. The you can equate the powers

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

Awesome! Thank you

OpenStudy (anonymous):

@campbell_st could you help me with one more? Or at least get me started? 5^2x-5=125